OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of three halves.

OpenStudy (anonymous):

help plz will give medal

OpenStudy (amistre64):

use the point slope for of a line

OpenStudy (anonymous):

hold on for answers

OpenStudy (amistre64):

given a point (a,b) and a slope (m) y-b = m(x-a)

OpenStudy (anonymous):

y - 5 = three halves(x + 1) y + 1 = three halves(x - 5) y + 5 = -three halves(x - 1) y - 1 = three halves(x + 5)